40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Mon Dec 28 12:35:57 2009, Jenne, Update, SUS, MC2 is having a bad day 
    Reply  Mon Dec 28 12:45:55 2009, Koji, Update, SUS, MC2 is having a bad day 
Message ID: 2458     Entry time: Mon Dec 28 12:45:55 2009     In reply to: 2457
Author: Koji 
Type: Update 
Category: SUS 
Subject: MC2 is having a bad day 

The MCL path of MC2 was in a strange state as the filters were activated as if it is in lock even though we had no lock. So I manually ran "mcdown". This reset the filters of the MCL path.


MC2 is having a bad day, and I'm not yet sure why.  It's to do with the damping though.  When the damping is off, after a little while it will settle to ~30mV or so on the Watchdog screen.  When I enable all of the outputs and then turn on the damping, the optic gets kicked up.  It's like there's a minus sign error somewhere, maybe in a bad burtrestore?  This has been going on since I did my morning bootfest.

It's started to sit down and play nicely now.  Is someone doing magic remotely that is fixing things that I hadn't figured out yet?


ELOG V3.1.3-